OpenStudy (anonymous):

Create your own two-step equation with NO SOLUTION as the solution. Show the steps to solving the equation and explain in words why your work proves that there are NO SOLUTIONS.

hero (hero):

x + 3 = x + 4 x - x = 4 - 3 0 = 1 False There is no solution because there exists no number, x, where the given equation produces an equal result.